OpenStudy (hba):

Which of the following satisfy f(x) >1 for all real values of x ? A) f(x)=x^2+2 B)f(x)=2sinx-2 C)f(x)=x^2-2 D)f(x)=|x+2|

OpenStudy (hba):

@AbhimanyuPudi

OpenStudy (anonymous):

i think A

OpenStudy (hba):

@gerryliyana Well My questions do not actually need answers :)

OpenStudy (hba):

Explanation Please.

OpenStudy (anonymous):

put whatever u want the value of x on f(x)=x^2+2, You will always get a f(x) >1

OpenStudy (anonymous):

A) x^2 is always >= 0 So, x^2+2 >= 2 If the values are greater than or equal to 2, then they are surely greater than 1, right?

OpenStudy (anonymous):

because there is quadratic that always positive and "+2" that always provide the value more than 1

OpenStudy (hba):

Ok Thanks Alot.

OpenStudy (anonymous):

welcome : )

OpenStudy (anonymous):

welcome :)